please could you help me solve for x:
ln(x-5)+ln(x-6)=ln(16-2x)

Answer by nerdybill(7384)

ln(x-5)+ln(x-6)=ln(16-2x)
ln(x-5)(x-6)=ln(16-2x)
(x-5)(x-6)=(16-2x)
x^2-11x+30 = 16-2x
x^2-9x+30 = 16
x^2-9x+14 = 0
(x-7)(x-2) = 0
x = {2, 7}
2 is an extraneous answer -- throw it out leaving:
x = 7